OpenStudy (anonymous):

how do you solve x^2+6x+9=1 by using the square root property?

myininaya (myininaya):

\[x^2+6x+9 =(x+3)^2 \text{ \right?}\]

myininaya (myininaya):

\[(x+3)^2=1 \]

myininaya (myininaya):

now take square root of both sides (x+3)=1 (x+3)=-1 x+3=1 x+3=-1

myininaya (myininaya):

x=1-3 x=-1-3

myininaya (myininaya):

x=-2 x=-4

OpenStudy (anonymous):

Thankyou!

OpenStudy (anonymous):

so for example in this problem, \[x ^{2}-8x+16=8\] it would be

OpenStudy (anonymous):

\[(x-4)^{2}=8\]

OpenStudy (anonymous):

right?

myininaya (myininaya):

yes so far good! :)

OpenStudy (anonymous):

and how do you get the last part again?

myininaya (myininaya):

you want to isolate x so first you need to isolate x-4 to do that you need to undo the square by square rooting both sides but that gives us two equations \[x-4=\pm \sqrt{8}\]

myininaya (myininaya):

\[x-4=\sqrt{8} \text{ or } x-4=-\sqrt{8}\]

OpenStudy (anonymous):

so you dont do the opposite ?

OpenStudy (anonymous):

no tin these problems?

myininaya (myininaya):

ok i'm going to do this problem and add one more step okay?

myininaya (myininaya):

before i do that answer me this do you know that \[|x|=\sqrt{x^2}\]

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

ohhh okay so the asnwer is qoinq to be as a square root?

myininaya (myininaya):

\[|x|=\pm x \text{ depending on if x is negative or positive }\]

OpenStudy (anonymous):

ohhhh !

myininaya (myininaya):

\[|-3|=-(-3)\] \[|3|=+3\]

myininaya (myininaya):

ok anyways lets do this problem one sec...

myininaya (myininaya):

\[(x-4)^2=8\] so this is the part you got to you understand this step am i correct?

OpenStudy (anonymous):

yes .

myininaya (myininaya):

\[\sqrt{(x-4)^2}=\sqrt{8}\] \[\text{ but } \sqrt{(x-4)^2}=|x-4|\] \[\text{ so we actually have } |x-4|=\sqrt{8}\]

myininaya (myininaya):

\[\text{ this means we have } x-4 = \pm \sqrt{8}\]

myininaya (myininaya):

what about this? is this good so far?

myininaya (myininaya):

take your time and reading it

OpenStudy (anonymous):

ohh so actually its square root of 8,

OpenStudy (anonymous):

so in \[x ^{2}-2x+1=2\]

OpenStudy (anonymous):

its \[(x-1)^{2}=2\]

OpenStudy (anonymous):

then it would be \[x-1=\pm \sqrt{2}\]

myininaya (myininaya):

yes yes

OpenStudy (anonymous):

i got it ? :D

myininaya (myininaya):

yep

myininaya (myininaya):

\[x=1 \pm \sqrt{2} \text{ after we add 1 on both sides}\]

OpenStudy (anonymous):

both sides ?

myininaya (myininaya):

of the equal sign

myininaya (myininaya):

we have x-1 on one side and plus or minus square root of 2 on the other side

OpenStudy (anonymous):

so it would be \[x=\pm \sqrt{2}+1\] ?

myininaya (myininaya):

you can write it like that if you want

myininaya (myininaya):

so on the previous one we didn't finish solving for x we had \[x-4=\pm \sqrt{8}\]

OpenStudy (anonymous):

then that one would be \[x=4\pm \sqrt{8}\]

myininaya (myininaya):

yes and some teachers may want you to ''simplify" that

myininaya (myininaya):

\[\sqrt{8}=\sqrt{4} \cdot \sqrt{2}=2 \sqrt{2}\]

myininaya (myininaya):

so \[x=4 \pm 2 \sqrt{2}\]

myininaya (myininaya):

good job itszel i'm happy to help people who want to learn

OpenStudy (anonymous):

so the answer would be \[x=4\pm2\sqrt{2}\] ?

myininaya (myininaya):

yes for \[x^2-8x+16=8\]

OpenStudy (anonymous):

thats the final answer?